How many four digit numbers greater than 6800 can be formed using the digits 3,5,6,7,8 where repetition of digits is allowed?

Answer by stanbon(75887)

How many four digit numbers greater than 6800 can be formed using the digits 3,5,6,7,8 where repetition of digits is allowed?
------
If thousands digit is 8 you get 1*5^3= 125 numbers
If thousands digit is 7 you get 1*5^3 = 125 numbers
If thousands digit is 6 the hundreds digit is 8 you get 1*1*5^2 = 25 numbers
------
Ans: 2*125 + 25 = 275 numbers
----------
Cheers,
Stan H.
--------------